OpenStudy (anonymous):

if the area of a trapezoid is 234 and b1 is x+4 and b2 is 2x-4 what is b1

OpenStudy (tylerd):

\[\frac{ b1+b2 }{ 2 } \times H = 234\]

OpenStudy (jdoe0001):

is there a "h" or height given for it?